# The Wisdom of a Crowd of Brains: A Universal Brain Encoder

### v1

@@ Abstract
Image-to-fMRI encoding is important for both neuroscience research and practical applications. However, such “Brain-Encoders” have been typically trained per-subject and per fMRI-dataset, thus restricted to very limited training data. In this paper we propose a Universal Brain-Encoder, which can be trained jointly on data from many different subjects/datasets/machines.
What makes this possible is our new voxel-centric Encoder architecture, which learns a unique “voxel-embedding” per brain-voxel. Our Encoder trains to predict the response of each brain-voxel on every image, by directly computing the cross-attention between the brain-voxel embedding and multi-level deep image features. This voxel-centric architecture allows the functional role of each brain-voxel to naturally emerge from the voxel-image cross-attention. We show the power of this approach to (i) combine data from multiple different subjects (a “Crowd of Brains”) to improve each individual brain-encoding, (ii) quick & effective Transfer-Learning across subjects, datasets, and machines (e.g., 3-Tesla, 7-Tesla), with few training examples, and (iii) use the learned voxel-embeddings as a powerful tool to explore brain functionality (e.g., what is encoded where in the brain).
$\star$\$ $\star$\$ footnotetext: Indicates equal contribution
@@ 1 Introduction
fMRI (functional MRI) has emerged as a powerful tool for measuring brain activity. This enables brain scientists to explore active brain areas during various functions and behaviors [1, 2, 3, 4, 5]. However, a human can spend only limited time inside an fMRI machine. This results in fMRI-datasets too small to span the huge space of brain functionality or visual stimuli (natural images). Moreover, the variability in brain structure and function responses between different people [6, 7, 8, 9] makes it difficult to combine data across individuals that have not been exposed to the same stimuli. All of these form severe limitations on current ability to analyze brain functionality.
@@
Figure 1: Overview. The Universal Image-to-fMRI Brain-Encoder can train jointly on multiple subjects & datasets. It leverages cross-attention between learned shared brain-voxel embeddings and deep image features.
@@
Image-to-fMRI encoding models, which predict fMRI responses to natural images, have greatly advanced the field. With the rise of deep learning, sophisticated encoding models have emerged  [10, 11, 12, 13, 14, 15], offering novel insights into brain function [16, 17, 18]. However, despite these advances, these models are primarily subject-specific and machine-specific, requiring extensive individual data (which is prohibitive) for effective training. This limits the practical use of existing brain-encoders, and prevents their ability to leverage cross-subject data. Attempts to create multi-subject encoders (e.g., [19, 20, 21, 22]) have so far been very restrictive (see Sec. 2). These approaches have thus far not demonstrated success in merging data from multiple fMRI datasets with different stimuli and varying acquisition settings (different machine resolutions, different scanning protocols, etc.)
@@
In this paper, we introduce the first-ever Universal Image-to-fMRI Brain-Encoder, which jointly trains on and integrates information from a collection of very different fMRI datasets acquired over the years (see Fig. 1). These multiple fMRI datasets provide multiple subjects exposed to very different image stimuli, scanned on different fMRI machines (3-Tesla, 7-Tesla), with varying number of brain-voxels. What makes this possible is our new Brain-Encoder architecture, which learns a unique “voxel-embedding” vector per brain-voxel (a small cube volume within the brain).
@@
Our Encoder trains to predict the fMRI response of each brain-voxel on each input image, by aggregating the cross-attention between the brain-voxel embedding and multi-level deep-features of the image. Other than the voxel-specific embedding (a learned 256-dimensional vector), all other network weights are shared across all subjects and voxels. This strategy, of learning meaningful brain-voxel embeddings via voxel-image cross-attention, provides several unique benefits: (i) The functional role of each brain-voxel naturally emerges. (ii) The Brain-Encoder architecture is not restricted to a predetermined number of voxels per fMRI scan, a common limitation among existing brain encoders. This allows to train the encoder jointly on subjects scanned using fMRI machines with different scanning resolutions. (iii) When a new subject/dataset is introduced, all that needs to be learned is the new subject’s voxels embeddings. Since this is captured by a small number of weights, it can be learned with few training examples.
The per-voxel Embedding puts a focus on individual voxel characteristics, independent of subject identity or fMRI dataset. This allows voxel functionality to be accurately captured across different subjects/datasets/machines. Moreover, the cross-attention mechanism between these voxel-embeddings and multi-level deep image features enables each brain-voxel to appropriately align with its corresponding “semantic level” (whether low or high).
We show the power of our approach to a variety of tasks, including: (i) Integrate information from many different fMRI datasets obtained by a “Crowd of Brains”. This wealth of training data gives rise to a Universal brain encoder, whose performance/accuracy significantly exceeds that of individually-trained (subject-specific) brain-encoders. (ii) Simple Transfer-Learning of the Universal encoder to new subjects and new datasets, with very few training data per subject. (iii) The learned voxel-embeddings provide a new tool to explore brain functionality, providing insights into what is encoded where in the brain.
@@
  • The first-ever Universal Brain-Encoder, which can successfully integrate data from multiple diverse fMRI datasets (old & new), collected on many different subjects, with different fMRI machines (3T, 7T), on very different image datasets.
  • Universal-encoding significantly improves over individually-trained subject-specific encoding.
  • Transfer-Learning of the Universal-Encoder to new subjects/datasets with few training data.
  • Learn functionally-meaningful brain-voxel embeddings via multi-level voxel-image cross-attention. This powerful representation further allows to explore brain functionality.
@@ 2 Related Work
Visual brain encoders, which map complex visual stimuli to brain activity, have significantly advanced over the years.Visual brain encoders have significantly advanced the field of neuroscience by mapping complex visual stimuli to brain activity. Initially, these models utilized linear regression between hand-crafted image features, to predict fMRI responses on images [23, 24]. Over time, the field has evolved to incorporate deep learning approaches both for image feature extraction and training [10, 11, 12, 13, 14, 15]. These models are typically subject-specific, due to substantial differences between brain responses of different people [6, 7, 8, 9]. This not only prevents broad generalization of these models, but also restricts their training due to the limited amout of data available per subject.
Multi-Subject fMRI Representations:
Traditionally, multi-subject fMRI studies have relied on anatomical alignment – i.e., canonical brain mapping to align all brains to a common anatomical space [25, 26, 27, 28]. This often results in poor functional correspondences (i.e., poor alignment of multi-subject fMRI responses), due to the varied nature of brain functionality across people [25, 29, 30, 31, 32]. More advanced methods for functional alignment were recently proposed, such as Hyper-Alignment [29, 33, 34, 35] or dimensionality reduction using auto encoders [36]. However, these approaches typically require shared data (i.e., same images seen by multiple subjects), which significantly restricts their applicability, and cannot be used across different (mutually-exclusive) fMRI datasets. Our method overcomes these limitations by introducing a unique "voxel-embedding" for each brain-voxel, generated through predicting its response to images (via multi-level voxel-image cross-attention). This approach allows for tailored functional learning and information sharing across different subjects/datasets, without the need for any shared data or same fMRI machine.
Multi-Subject Brain Encoders:
Few attempts have been made to develop models that can benefit from multi-subject brain data [19, 20, 21, 22]. None of these approaches, however, have demonstrated the capability to integrate data across multiple subjects from different datasets, with varying fMRI resolutions, or in the absence of shared data. Methods like [19] use Multi-Subject fMRI representations to integrate data from different subjects but, as mentioned above, these require some shared-data. Other approaches employ end-to-end multi-subject encoders with partially shared weights [20], or use one subject’s encoder parameters as the basis for another [21], or leverage pre-trained encoder outputs for new subject adaptation [22]. None of those approaches have demonstrated effectiveness on diverse datasets and machines. Moreover, they usually treat each fMRI scan as a single whole entity, thus preventing effective learning and sharing of functional knowledge across different brain-voxels (whether of the same person or different people). In contrast, in our voxel-centric Universal Brain-Encoder all network weights are shared across all brain-voxels (for all subjects/datasets/machines), except for the unique voxel-specific embedding (learned via voxel-image cross-attention). This allows our model to learn shared voxel-functionality across different subjects, datasets, and fMRI machines.
@@ 3 The Universal-Encoder
Our Universal Encoder facilitates joint training on data from multiple subjects across various fMRI datasets, where subjects were exposed to completely different image stimuli and scanned using fMRI machines with differing resolutions (see Fig. 1). We first provide an overview of the method, followed by a detailed explanation of the architecture and training process.
Figure 2: Universal-Encoder Architecture. Input: An image & a Voxel-ID; Output: The predicted voxel activation on that image. The model has 3 main components: (a) Feature Extraction Block – extracts multi-scale image features; (b) Learned Voxel-Embedding – captures the unique functionality of each voxel; (c) Cross-Attention Block – establishes the connection between voxel-functionality and relevant visual information.
@@ 3.1 Overview of the Approach
The core innovation of our encoder lies in its integration of brain data with image features through a brain-image cross-attention mechanism. For effective integration, we developed a per-voxel learned embedding (for each brain-voxel of each subject). Each voxel-embedding is randomly initialized, and is refined during training to accurately capture the functionality of the voxel. Given an image and a voxel embedding, the encoder extracts various image features using a DINO-v2 [37] adapted model as our feature extraction block. It outputs image features from different intermediate layers of DINO, allowing each brain-voxel to attend to the appropriate semantic levels of features that align with its functionality. These image features, along with a specific voxel embedding, are processed through the cross-attention block to integrate them effectively and predict the voxel’s activation in response to the given image. Note that this architecture is indifferent to the number of brain-voxels in each fMRI scan, hence is applicable to any fMRI data.
@@
Our training process optimizes all 3 components simultaneously – the voxel embedding, the feature extraction block, and the cross-attention block – with the common goal of predicting the voxel response to the input image. This joint learning framework develops meaningful voxel embeddings, that not only improve voxel response prediction, but also implicitly captures its functional role in the brain. Our encoder and all associated weights are shared across all brain-voxels (for all subjects/datasets/machines), differing only in the per-voxel embeddings. This design ensures that each brain-voxel embedding is determined by its functional characteristics, rather than by its physical location in the brain or the subject’s identity. The shared voxel embedding space supports integration of information across different voxels (whether within a single brain or across different brains). Importantly, our approach does not require subjects to have viewed the same images, nor to have been scanned in the same machine. This allows, for the first time, to integrate information from numerous fMRI datasets, collected by different groups across the globe over many years! We refer to this as the “Wisdom of a Crowd of Brains”.
@@
Our proposed Universal-Encoder provides a powerful means for integrating data from multiple subjects, both within and across different fMRI datasets. These are empirically evaluated in Sec. 4. Moreover, it learns the functional role of each brain voxel, and maps functionally-similar brain voxels (both within the same brain, and across different brains) to nearby Voxel-Embeddings. This provides a powerful tool to explore the human brain and discover new functional regions within it. This is demonstrated in Sec. 5. What facilitates such advanced brain exploration is the enormous number of images that the large “crowd of brains” has collectively seen (which is prohibitive for a single subject).
@@ 3.2 Architecture and Training
The architecture of our encoder is designed to receive 2 inputs: (i) an image, (ii) a Voxel-ID, and outputs a single scalar value – the predicted fMRI voxel activation on that image. The encoder architecture comprises three main components: the image features extraction block (Fig. 2a), a shared voxel embedding space (Fig. 2b), and the Voxel-Image cross-attention block (Fig. 2c).
Image Features Extraction Block
@@
(Fig. 2a). This block utilizes an adapted DINO v2 model [37] to derive multi-scale image features. Features are extracted from L=5 intermediate layers of the DINO V2 VIT-G/14 model (layers 1,6,12,18,24), where lower layers capture low-level image features and higher layers provide more semantic information. This hierarchical feature extraction is crucial, as voxels in the visual cortex correspond to a range of image attributes – from simple visual details to complex semantic content. Each layer’s features are projected to a lower-dimension C (using a linear layer), and are then concatenated along another dimension of length L (the number of layers). Given that DINO operates on P image patches, the final feature output is of size L×P×C. In order to transform Dino features into features that are suitable for predicting brain activity, we used a LoRA inspired approach [38], that is more suitable for data-limited settings (see A.2 for details).
Per-Voxel Embedding
(Fig. 2b). Each brain voxel of each subject is assigned with a vector of length E=256. This E-dimensional vector (“Embedding”) is initialized randomly, and is learned during training to maximize the prediction of the voxel fMRI activation on images (computed via cross-attention of the voxel-embedding with image features). This voxel embedding is designed to capture the unique functionality of each voxel, facilitating the mapping of brain voxels from different subjects to the same functional space. Note that the final voxel-embedding is not image-dependant.
Cross-Attention Block
@@
(Fig. 2c). The voxel-image cross-attention block establishes the connection between voxel functionality and relevant visual information. This block includes three sequential components: (i) Spatial-attention, (ii) MLPs (multi-layer perceptrons), and (iii) Functional-attention. The Spatial-attention component allows the voxel embedding to select relevant locations within the image, while the Functional-attention component selects the relevant features at these locations. Both components are essential, as different brain voxels have varying image receptive fields (in both location and size – some have a small receptive field within the image, others are affected by the entire image), and different functionalities (e.g., low-level versus high-level semantic features). Given the features of the input image (referred to as “input features” from here on), the Spatial attention enables each voxel to focus on its corresponding spatial location within the image, effectively selecting features from the appropriate image patches. Using attention notations, the “Values” V are the input features, with dimension L×P×C. The “Query” vector $\mathbb{q}$ is the Voxel-Embedding, transformed by a linear layer which preserves its size (1×E). The "Keys" K are derived by adding a learned per-patch positional encoding (size P×E) to the input features projected to the embedding size E. The output is calculated by $softmax(\mathbb{q}{K_{L}}^{T})V_{L}$ for each of the L layers separately (a weighted summation across the spatial dimension P), outputting vectors of size L×C. The spatially averaged features are then fed to MLPs (a separate 2-layered MLP for each of the L image-feature layers), maintaining the dimensions L×C. Lastly, the Functional-attention performs a weighted summation of the spatially-attended features to derive a single scalar voxel activation. In this layer, $\mathbb{v}$ represents the flattened MLP output (size 1x(L*C) ), $\mathbb{q}$ is the voxel embedding itself, and K is learned feature embedding that has an entry for each of the LXC features (size (L*C)XE). The output is calculated via $(\mathbb{q}K^{T})\mathbb{v}^{T}$. This block outputs the voxel prediction as a scalar value.
@@
The model is trained end-to-end with all components learned together, where the objective of the model is to correctly predict the voxel activation on each input image. Our training dataset comprises images along with corresponding fMRI scans, collected on many different subjects from multiple different fMRI datasets. Training batches are constructed from 32 randomly selected images, where for each image we randomly sample 5000 voxel indices along with their corresponding fMRI activations for prediction. Each subject (brain) has its own unique voxel indices. The model is trained using the Adam optimizer with a learning rate of 1e-3. For the loss function, we employ the same loss as in [14], using an affine combination of MSE loss and cosine proximity:
$\mathcal{L}\left(\hat{r},r\right)=\alpha\cdot\text{MSE}\left(\hat{r},r\right)-\left(1-\alpha\right)\cos\left(\angle\left(\hat{r},r\right)\right)$ (1)
where $\hat{r}$ and $r$ are the predicted and measured fMRI activations, respectively, and $\alpha=0.1$. Training the Universal-Encoder jointly on 8 NSD subjects (see ‘Datasets” below), takes $\sim$1 day on a single Quadro RTX 8000 GPU. Inference time (Image-to-fMRI encoding) takes $\sim$50 ms per-image.
Figure 3: Qualitative Evaluation of the Universal-Encoder. (a) Visual comparison of Real vs. Encoder-predicted fMRIs for 3 test images. (b) Top 5 retrieved images for each "Query" test-fMRI (see text for details).

### v2

@@ Abstract
Image-to-fMRI encoding is important for both neuroscience research and practical applications. However, such “Brain-Encoders” have been typically trained per-subject and per fMRI-dataset, thus restricted to very limited training data. In this paper we propose a Universal Brain-Encoder, which can be trained jointly on data from many different subjects/datasets/machines. What makes this possible is our new voxel-centric Encoder architecture, which learns a unique “voxel-embedding” per brain-voxel. Our Encoder trains to predict the response of each brain-voxel on every image, by directly computing the cross-attention between the brain-voxel embedding and multi-level deep image features. This voxel-centric architecture allows the functional role of each brain-voxel to naturally emerge from the voxel-image cross-attention. We show the power of this approach to: (i) combine data from multiple different subjects (a “Crowd of Brains”) to improve each individual brain-encoding, (ii) quick & effective Transfer-Learning across subjects, datasets, and machines (e.g., 3-Tesla, 7-Tesla), with few training examples, and (iii) we show the potential power of the learned voxel-embeddings to explore brain functionality (e.g., what is encoded where in the brain).
@@ 1 Introduction
fMRI (functional MRI) has emerged as a powerful tool for measuring brain activity. This enables brain scientists to explore active brain areas during various functions and behaviors (Kanwisher et al., 1997; Epstein & Kanwisher, 1998; Downing et al., 2001; Tang et al., 2017; Heeger & Ress, 2002). However, a human can spend only limited time inside an fMRI machine. This results in fMRI-datasets too small to span the huge space of brain functionality or visual stimuli (natural images). Moreover, the variability in brain structure and function responses between different people (Riddle & Purves, 1995; Frost & Goebel, 2012; Conroy et al., 2013; Zhen et al., 2015) makes it difficult to combine data across individuals that have not been exposed to the same stimuli. All of these form severe limitations on current ability to analyze brain functionality.
@@
Figure 1: Overview. The Universal Image-to-fMRI Brain-Encoder trains jointly on multiple subjects & datasets. It learns to predict fMRI activation of each brain-voxel on any image, via cross-attention between learned brain-voxel embeddings and deep image features.
@@
Image-to-fMRI encoding models, which predict fMRI responses to natural images, have greatly advanced the field. With the rise of deep learning, sophisticated encoding models have emerged  (Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018c, a; Beliy et al., 2019; Gaziv et al., 2022), offering novel insights into brain function (Tang et al., 2024; Henderson et al., 2023; Gu et al., 2023; Luo et al., 2024, 2023). However, these models are primarily subject-specific and machine-specific, requiring extensive individual data (which is prohibitive) for effective training. This limits the practical use of existing brain-encoders, and prevents their ability to leverage cross-subject data. Attempts to create multi-subject encoders (e.g., (Van Uden et al., 2018; Khosla et al., 2020; Wen et al., 2018b; Gu et al., 2022)) have so far been very restrictive (see Sec. 2), showing no generalization across different datasets/machines.
@@
In this paper, we introduce the first-ever Universal Image-to-fMRI Brain-Encoder, which jointly trains and integrates information from a collection of very different fMRI datasets acquired over the years (see Fig. 1). These multiple fMRI datasets provide multiple subjects exposed to very different image stimuli, scanned on different fMRI machines (3-Tesla, 7-Tesla), with varying number of brain-voxels (a “brain-voxel” is a tiny 3D cube of brain). What makes this possible is our new brain-voxel centric approach, which captures subject-specific and voxel-specific functionality, through a learned “voxel-embedding” vector. During training, each voxel-embedding learns to encode the unique visual functionality of the corresponding brain-voxel, while still leveraging shared patterns across different brain-voxels and different subjects (via all other shared network components). Our voxel-centric approach is in contrast to all previous fMRI-centric methods, which treat each fMRI scan as a single complete entity. They can thus exploit only shared information of the entire fMRI across different scans, overlooking the frequent similarity between different voxels within a single fMRI scan.
Our approach builds on the observation that similarity between individual voxels – both within & across brains, is more likely to occur than similarity between entire fMRI scans across different brains.
@@
Our Encoder trains to predict the fMRI response of each brain-voxel on any input image, by aggregating the cross-attention between the brain-voxel embeddings and multi-level deep-features of the image. Each brain voxel (of every subject) has a corresponding voxel embedding. This embedding is merely a vector of size 256, which learns to capture what this brain-voxel is sensitive to: whether it attends to low-level image features or to high-level ones; whether it cares about the position of the feature in the image or not; etc. This voxel-embedding is initialized randomly, and is optimized end-to-end during training. Other than the voxel-specific embedding vectors, all other network weights are shared across all voxels of all subjects. The per-voxel Embedding puts a focus on individual voxel characteristics, independent of subject identity or fMRI dataset. This strategy provides several unique benefits: (i) The functional role of each brain-voxel naturally emerges. (ii) Being voxel-centric, the Universal Brain-Encoder architecture is indifferent to the number of voxels per fMRI scan, which enables joint training on data from fMRI machines of different scanning resolutions (e.g., 3T, 4T, 7T). (iii) Once trained, transfer learning to a new subject or a new fMRI dataset/machine requires only few new training examples.
@@
  • The first-ever Universal Brain-Encoder, successfully integrates data from multiple diverse fMRI datasets (old & new), from many different subjects, and different fMRI machines (3T, 4T, 7T). This is made possible by our voxel-centric encoding model.
  • Combining data from “Crowds of Brains/Datasets” provides significantly more training data. This leads to a significant encoding improvement compared to any individually-trained subject-specific encoder.
  • Transfer-Learning to new subjects/datasets/machines is obtained with very few training data.
  • The learned “voxel-embeddings” provide a new tool to explore brain functionality, providing insights into what is encoded where in the brain. Such advanced brain exploration is facilitated by the enormous number of images collectively seen by the “crowd of brains/datasets” (which is prohibitive for any single subject).
@@ 2 Related Work
Methods for Mapping visual stimuli to brain activity (“Image-to-fMRI Encoding”), have significantly advanced over the years, with contributions to Neuroscience. Initially, these models utilized linear regression between hand-crafted image features, to predict fMRI responses on images (Kay et al., 2008; Naselaris et al., 2011). Over time, the field has evolved to incorporate deep learning approaches both for image feature extraction and training (Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018c, a; Beliy et al., 2019; Gaziv et al., 2022; Wang et al., 2023; Takagi & Nishimoto, 2023). These models are typically subject-specific, due to substantial differences between brain responses of different people (Riddle & Purves, 1995; Frost & Goebel, 2012; Conroy et al., 2013; Zhen et al., 2015). This restricts the encoder to subject-specific data, and limits its generalization across subjects.
Multi-Subject Brain Models: Efficiently integrating data from multiple subjects is challenging due to anatomical and functional differences between brains. While anatomical alignment (Mazziotta et al., 2001; Talairach, 1988; Fischl, 2012; Dale et al., 1999) aligns brains to a common anatomical space, it fails to provide accurate functional correspondences(Mazziotta et al., 2001; Haxby et al., 2011; Yamada et al., 2015; Brett et al., 2002; Wasserman et al., 2024). Although we deal here with Multi-Subject Brain Encoding, we review both Multi-Subject Encoding & Decoding.
Decoders:
In fMRI-to-Image decoding, the goal is to reconstruct an image given its corresponding fMRI brain recording. Recent advanced fMRI decoders leverage multi-subject data (e.g., ‘MindBridge’ (Wang et al., 2024), ‘MindEye2’ (Scotti et al., 2024b), (Gong et al., 2024) and (Han et al., 2024)), by projecting the subject-specific fMRIs of all subjects into a shared space, which is then fed into a shared image-decoding network. However, the problems of multi-subject fMRI Decoding & Encoding are fundamentally different. In multi-subject Decoding, the network inputs reside in different spaces (individual subject-specific fMRIs), whereas the network outputs reside in a shared space (the Image-Space). Thus projecting to a shared space before decoding is a valid solution. In contrast, in multi-subject Encoding, the inputs (images) reside in a shared space, whereas the outputs (individual subject-specific fMRIs) reside in completely different spaces. Thus the above approach of ‘MindBridge’&‘MindEye2’ will not apply to multi-subject Encoding, as projecting to a shared space will lose most subject-specific fMRI variability (output variability). This is resolved by our new voxel-centric model.
Encoders:
Few attempts have been made to leverage multi-subject data for Image-to-fMRI Encoding. Van Uden et al. (2018) integrated data from multiple subjects, but required them to be exposed to same images, which is very restrictive. Our approach requires NO shared data between subjects. Wen et al. (2018b) and Gu et al. (2022) finetune models of one subject to another. This allows per subject adaption, but largely limits the information sharing as they are not trained jointly. Lastly, Khosla et al. (2020) use a model with part shared and part subject-specific. However, to obtain good results they predict only a small subset of voxels which have high correlation across brains.
Finally, all previous methods (both Encoding and Decoding) are predominantly fMRI-centric. Namely, they treat each fMRI scan as a single entity, relying on shared embeddings of entire fMRI scans across subjects. Thus they can only exploit shared representations of entire fMRIs across different scans, overlooking the frequent similarity between individual brain-voxels within a single fMRI scan (a single brain), as well as across different brains. In contrast, our Voxel-centric model fundamentally shifts away from this paradigm by sharing network weights across all brain voxels – both within and across brains. Since similarity between individual voxels (within and across brains) is more frequent than similarity of entire fMRI scans, our model can effectively integrate multi-subject data.
@@ 3 The Universal-Encoder
Our Universal Encoder facilitates joint training on multi-subject data from diverse fMRI datasets, with subjects exposed to completely different image stimuli and scanned on fMRI machines with very differing resolutions (see Fig. 1).
Figure 2: Universal-Encoder Architecture. Input: an image & a brain-voxel index (a pointer to its Voxel-Embedding vector); Output: The predicted fMRI activation of this brain-voxel on that image. The model has 3 main components: (a) Feature Extraction Block – extracts multi-scale (DINO-adapted) image features; (b) Learned Voxel-Embedding – captures the unique functionality of each voxel; (c) Cross-Attention Block – establishes the connection between voxel-functionality and relevant image features
@@ 3.1 Overview of the Approach
The core innovation of our encoder lies in the voxel-centric architecture, as well as in the brain-image cross-attention mechanism. Each brain-voxel of each subject is assigned a unique corresponding “voxel-embedding” (a vector of size 256 – Fig. 2b). This embedding vector is initialized randomly and is optimized during training to learn to predict the fMRI response of that brain-voxel on any image (via our voxel-image cross-attention mechanism).
Since this embedding is voxel-specific (not image-specific), yet learns to predict this voxel’s activation on any image, it must therefore encode inside its vector the “functionality” of this specific brain-voxel (i.e., what it is sensitive to in visual data): whether it attends to low-level image features or to high-level semantic ones; whether it cares about the position of the feature within the image or not; etc.
@@
The shared network components (shared across all brain-voxels of all subjects) include the feature extraction block (Fig2a), and the voxel-image cross-attention block (Fig.2c). Given any image, our encoder extracts from it various image features (from low-level to high-level) using a DINO-v2 (Oquab et al., 2023) adapted model as our feature extraction block. It outputs image features from different intermediate layers of DINO, allowing each brain-voxel to attend to the appropriate semantic levels of features that align with its functionality. These image features, along with a specific voxel-embedding, are processed through the cross-attention block, to effectively integrate and predict the voxel’s activation in response to the given image. Note that this architecture is indifferent to the number of brain-voxels in each fMRI scan, hence applies to any fMRI scan.
@@
Our training process optimizes all 3 components simultaneously – the voxel embedding, the feature extraction block, and the cross-attention block – with the sole goal of predicting the voxel fMRI response to the input image. This joint learning framework develops meaningful voxel embeddings, that not only improve voxel response prediction, but also implicitly ‘encodes’ its functional role in the brain. Our encoder and all associated weights are shared across all brain-voxels (for all subjects/datasets/machines), differing only in the per-voxel embeddings. This design ensures that each brain-voxel embedding is determined by its functional role, rather than by its physical location in the brain or the subject’s identity. The shared voxel embedding space supports integration of information across different voxels (whether within a single brain or across different brains). Importantly, our approach does not require subjects to view the same images, nor to be scanned on the same machine. This allows, for the first time, to integrate information from numerous fMRI datasets, collected by different groups around the globe over many years! We refer to this as the “Wisdom of a Crowd of Brains”.
@@
The ability of our Universal-Encoder to integrate data from multiple subjects within/across fMRI datasets is empirically evaluated in Sec. 4. We further demonstrate the ability of our Encoder to learn the functional role of each brain-voxel, and map functionally-similar brain-voxels (within the same brain, and across different brains) to nearby Voxel-Embeddings. This provides a powerful tool to explore the human brain on huge amounts of data collectively acquired by many brains, thus discover new functional brain regions. This is demonstrated in Sec. 5.
@@ 3.2 Architecture and Training
Our encoder architecture receives 2 inputs:  (i) an image, (ii) a brain-voxel index (which is merely a pointer to this brain-voxel’s Embedding vector). It outputs a single scalar value – the predicted fMRI response of this voxel on that image. The encoder architecture comprises 3 main components: (i) the shared image features extraction block (Fig.2a), (ii) the voxel embedding vectors (Fig.2b), (iii) the shared Voxel-Image cross-attention block (Fig.2c).
Image Features Extraction Block:
@@
(Fig. 2a). This block utilizes an adapted DINO model (Oquab et al., 2023) to derive multi-scale image features. Features are extracted from L=5 intermediate layers of the DINO-v2 VIT-L/14 model (layers 1,6,12,18,24), where lower layers capture low-level image features and higher layers provide more semantic information. This hierarchical feature extraction is crucial, as voxels in the visual cortex correspond to a range of image attributes – from simple visual details to complex semantic content. Each layer’s features are projected to a lower-dimension C (via a linear layer), and are then concatenated along another dimension of length L (the number of layers). Since DINO operates on P image patches, the final feature output is of size L×P×C. In order to transform Dino features into features suitable for predicting brain activity, we used a LoRA inspired approach (Hu et al., 2021), that is more suitable for data-limited settings (see A.2 for details).
Figure 3: Qualitative Evaluation of the Universal-Encoder. (a) Visual comparison of Real vs. Encoder-predicted fMRIs for 3 test images. (b) Top 5 retrieved images for each ”Query” test-fMRI.
(See text for more details)
Per-Voxel Embedding:
(Fig. 2b). Each brain voxel of each subject is assigned a voxel-specific vector of length E=256. This E-dimensional vector (“Embedding”) is initialized randomly, and is optimized during training to accurately predict this voxel’s fMRI activation on different images (predicted from the cross-attention between the voxel-embedding and the image features). Since this embedding is voxel-specific (not image-specific), it must learn to encode inside its vector the “functionality” of this specific brain-voxel (i.e., what it cares about in images). Note that: (i) While each optimized embedding is voxel-specific, the remaining network components are shared by all voxels of all subjects. This allows to train all the shared components of our Encoder on data from multiple subjects/datasets/machines, even if they have significantly different numbers of brain voxels. (ii) Such joint training further facilitates the mapping of brain voxels from different brains of different subjects to the same embedding space. This allows for shared functional regions across different brains (who have never seen any shared data), to naturally surface out and be discovered. This is discussed in Sec. 5, with a few such examples shown in Fig. 7. (iii) Note that unlike the common use of the term “embedding” in Deep-Learning, our voxel-embeddings are not an output of any sub-network. These embedding vectors are initialized randomly, and are optimized individually during training along with all other shared network components.
Cross-Attention Block:
(Fig. 2c). The voxel-image cross-attention block establishes the connection between voxel functionality and relevant visual information. This block includes 3 components: (i) Spatial-attention, (ii) MLPs, and (iii) Functional-attention. The Spatial-attention component allows the voxel embedding to select relevant locations within the image, while the Functional-attention component selects the relevant features at these locations. Both components are essential, as different brain voxels have varying image receptive fields (some voxels have small, localized receptive fields, while others are influenced by the entire image), and different functionalities (e.g., low-level versus high-level semantic features).
More specifically:
@@
(i) Given the features of the input image (referred to as “input features” from here on), the Spatial attention enables each voxel to focus on its corresponding spatial location within the image, effectively selecting features from the appropriate image patches. Using attention notations, the “Values” V are the input features, with dimension L×P×C. The “Query” vector $\mathbb{q}$ is the Voxel-Embedding, transformed by a linear layer which preserves its size (1×E). The ”Keys” K are derived by adding a learned per-patch positional-embedding (size P×E) to the input features projected to the embedding size E. The output is calculated by $softmax(\mathbb{q}{K_{L}}^{T})V_{L}$ for each of the L layers separately (a weighted summation across the spatial dimension P), outputting vectors of size L×C.
(ii) The spatially averaged features are then fed to MLPs (a separate 2-layered MLP for each of the L image-feature layers), maintaining the dimensions L×C.
(iii) Lastly, the Functional-attention performs a weighted summation of the spatially-attended features to derive a single scalar voxel activation. In this layer, $\mathbb{v}$ represents the flattened MLP output (size 1x(L*C) ), $\mathbb{q}$ is the voxel embedding itself, and K is learned functional-embedding that has an entry for each of the LXC features (size (L*C)$\times$×E). The output is calculated via $(\mathbb{q}K^{T})\mathbb{v}^{T}$. This block outputs the voxel prediction as a scalar value.
@@
The model is trained end-to-end to correctly predict the voxel fMRI activation on each input image. The loss is: $\mathcal{L}\left(\hat{r},r\right)=\alpha\cdot\text{MSE}\left(\hat{r},r\right)-\left(1-\alpha\right)\cos\left(\angle\left(\hat{r},r\right)\right),$ where $\hat{r}$ and $r$ are the predicted and measured fMRI activations, respectively, and $\alpha$=0.1. Training the Universal-Encoder jointly on 8 NSD subjects (see “Datasets” below), takes $\sim$1 day on a single Quadro RTX 8000 GPU. Inference time (full Image-to-fMRI encoding) takes $\sim$50 ms per-image. See further training details in appendix A.3.